A recipe calls for 2c of sugar to 5c of flour. How much flour would be added to 5c of sugar?
Found 2 solutions by stanbon, roseo:
Answer by stanbon(75887)

A recipe calls for 2c of sugar to 5c of flour. How much flour would be added to 5c of sugar?
-----------------
Proportion:
x/5c = 5c/2c
x = 25c/2c
x = 12.5c (amt. of flour needed with 5c of sugar)
=====================================================
Cheers,
Stan H.

Answer by roseo(33)

Think of this as a proportion problem: 2 cups of sugar to 5 cups of flour and then 5 cups for sugar to how much flour?
2/5 = 5/x
Use cross multiplication for 2x = 25 then x = 12.5 or 12.5 cups of flour